# Effective synchronization amid noise-induced chaos

### v2

@@ Abstract
Two remote agents with synchronized clocks may use them to act in concert and communicate. This necessitates some means of creating and maintaining synchrony. One method, not requiring any direct interaction between the agents, is to expose them to a common, environmental, stochastic forcing. This “noise-induced synchronization” only occurs under sufficiently mild forcing; stronger forcing disrupts synchronization. We investigate the regime of strong noise, where the clocks’ relative phases evolve chaotically. Using a simple realization of disruptive noise, we demonstrate effective synchronization. First, although the relative phases of the two clocks varied erratically, we confirm that they became statistically independent of initial conditions and hence equivalent after a well-defined timescale. Second, we show that an agent can estimate an effective phase that closely agrees with the other’s phase. Thus, synchronization is practically attainable beyond the regime of conventional noise-induced synchronization. We finally discuss how it might be used in living systems.
Significance
In ecology and neurobiology, independent living agents may anticipate one another’s behavior by means of a shared measure of time, e.g., simple chemical clocks requiring constant synchronization. We consider the effect of random events in the agents’ common environment that disrupt synchronization. We show that despite the resultant chaotic dynamics of their clocks, the two agents can independently analyze the impact of the recent environmental noise on their clock phases to recover effective synchronization. This scheme offers a route to explain how coordinated behavior can emerge amid shared disruptive noise. An example is the puzzling large information transfer encoded in spike timings of sensory neurons.
@@ I Introduction
One such means is noise-induced synchronization, occurring when identical noninteracting oscillators are subjected to a common, nonlinear, stochastic perturbation from their environment [12, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. It is distinct from synchronization due to interactions between the clocks and from phase locking unto an external clock. Instead, synchronization is enabled via the clocks’ shared ambient noise. As such, noise-induced synchronization is the only means of maintaining synchronization between noninteracting clocks in a generic environment. When the noise is sufficiently mild, it can bring the two oscillators into synchrony over time. This has been well characterized experimentally in a range of physical contexts, such as electric circuits [21], quantum optics signaling [25], and even gene regulatory networks [26]. Synchronization can enable time-based communication between a sender and a receiver. If the sender emits an impulse when its clock is at a particular phase, the receiver can infer this phase by noting its own clock’s phase at the moment of the impulse.
@@
There is no guarantee that two oscillators will synchronize under a given noise. In general, synchronization requires the strength of the noise to lie below a threshold. Beyond this threshold, the relative phase of the two oscillators devolves into chaos [18, 24]. In the latter case, synchronization is lost and with it the means of time-based communication outlined above. Nevertheless, we show below that a form of synchronization persists in the statistical properties of these erratically evolving phases. It permits each agent to define an effective phase that agrees with that of other agents to a quantitatively-predictable extent, thus enabling them to behave as though synchronized.
@@
We demonstrate this behavior using a simplified type of noise, consisting of randomly timed impulses. For example, this may represent stochastically timed electric shocks delivered to a collection of firing neurons [7, 4, 6] or injections of a chemical affecting the cell cycle into a cell culture [3]. We study the effect of these perturbations on a distribution of phases. Successive impulses create a sequence of phase distributions that in general depend on the initial distribution and the timings of the impulses. In below-threshold noise-induced synchronization, these distributions become narrow and unimodal, concentrated at a single phase. For noises above the threshold, the distributions remain spread over the entire phase circle [24].
@@
The above-threshold effective synchronization we report here arises by virtue of two properties: First, though the distributions change markedly with each impulse, we find that they become independent of the initial state. This means that two initially different oscillator ensembles become statistically equivalent when compared at a given instant. Second, the distributions can be tightly bunched into a few sharp peaks. Indeed, when the threshold of synchronization is approached from above, the distributions show a high degree of order, since their average entropy becomes arbitrarily negative [24].
@@ II.1 Impulsive perturbations as phase maps
The effect of a kick on the clock can be compactly expressed via a phase-reduction prescription [22]: The oscillator traces a periodic loop in its dynamical manifold with a period $T$. We may label the points on this loop by a “phase position” $\varphi$ ranging from 0 to $1$, with a designated point assigned as the phase origin $\varphi\equiv 0$. Other points are assigned phase values equal to the time $\Delta t$ required for the oscillator to reach them from the phase origin: $\varphi\equiv\Delta t/T$. This construction is equivalent to finding the action-angle variables in Hamiltonian mechanics [27]. Thus, the phase of an oscillator that started at $\varphi^{(0)}$ advances at a constant speed, $\varphi^{(0)}+t/T$, where $\mathrm{mod}\,1$ is implied in any algebraic operation over the phases.
@@
A unique phase map $\psi(\varphi)$ can be determined for a given pairing of a dynamical system and a specified kick, as illustrated in SI Appendix, Sec. S2 [28, 29]. In biological contexts, $\psi(\varphi)-\varphi$ is called the phase response curve, quantifying the response of circadian clocks [2, 30] and neural networks [31] to environmental changes. In these cases, the clocks are a living organism and firing neurons, and the perturbations can be sporadic light impulses or shot noise from sensory neurons.
We study the case where this deterministic perturbation $\psi(\varphi)$ is applied at random times. [^1] We assume that the waiting times in between kicks are sufficiently long so the oscillators have relaxed back to the limit cycle. For the case of possibly insufficiently-long times, see Ref. [32]. For simplicity, we may capture the effect of a $k$’th kick by observing the oscillators at an integer number of cycles after the latest kick. The unkicked position of the phase origin has continued its periodic motion and thus returns to the phase origin at such a moment. Likewise, the same phase is observed for the kicked oscillator after each completed cycle; we denote this phase as $\varphi^{(k)}$. Thus, the phase-map shift $\psi(\varphi^{\prime})$ from the unkicked phase origin is simply $\varphi^{(k)}$. To complete the iteration, we must express the phase position at the moment of the $k$th kick $\varphi^{\prime}$ in terms of the previous kick’s phase $\varphi^{(k-1)}$. During the random waiting time between the k1k-1’st and the $k$’th kick, the oscillator has gained a phase that we denote by $\beta^{(k)}$. Thus, $\varphi^{\prime}=\varphi^{(k-1)}+\beta^{(k)}$. Combining, we infer the effect of a single iteration,
@@
Due to the $\mathrm{mod}\,1$ constraint, only the fractional part of the waiting time in between kicks matters, which we denoted $\beta^{(k)}$. For simplicity, we assume that $\{\beta^{(1)},\beta^{(2)},\ldots\}$ are identical independent uniformly-distributed random phases, $\beta^{(k)}\in[0,1)$. Equation (1), therefore, amounts to a Markov dynamics in the discrete “time” $k$, controlled by the phase map $\psi(\varphi)$ and the drawn $\beta^{(k)}$’s. Below, we compare the fates of two ensembles of oscillators exposed to the same sequence of $\beta^{(k)}$’s. [^2] We consider synchronization of distinct initial phase distributions under the influence of a shared forcing, as opposed to identical initial phase distribution evolving under different forcings. We average over $\{\beta^{(k)}\}$’s only after computing differences between the two agents obtained with fixed $\{\beta^{(k)}\}$’s.
@@
where the “gain parameter” $A$ determines the strength of the perturbations; when $A=0$, $\psi(\varphi)=\varphi$, so the kick has no effect. We plot this $\psi(\varphi)$ for $A=9.32$ in SI Appendix, Fig. S4, solid red curve. For completeness, we have repeated the paper’s quantitative analyses also for the quintic phase map, $\psi(\varphi)=\varphi+A^{\prime}\varphi(1-\varphi)(\Psi_{1}-\varphi)(\Psi_{2}-\varphi)(C-\varphi)$, where $C=(1-\Psi_{1})(1-\Psi_{2})/[\Psi_{1}\Psi_{2}+(1-\Psi_{1})(1-\Psi_{2})]$ is such that $(d\psi/d\varphi)|_{\varphi=1}=(d\psi/d\varphi)|_{\varphi=0}$. We will show the results for the cubic phase maps; we observed that the same effective synchronization occurs for the quintic ones.
@@ II.2 Lyapunov exponent and synchronization
With maps of this form, strong perturbations will typically be characterized by an increased gain parameter $A$. Once perturbed, the quantity $|d\psi/d\varphi|$ measures the degree of spreading of a local interval around $\varphi$. This “spreading factor” therefore is closely related to the (in)ability of the phase map to synchronize adjacent phases. Hence, a measure of effective strength of the kicks can be formulated by averaging this spreading factor. Accordingly, we define $\Lambda$ — the average of the log-spreading factor, [^3] The integration measure is uniform since the waiting times are distributed uniformly $\beta\in[0,1)$. Therefore, every oscillator is likely to experience a kick with equal probability everywhere along the phase circle.
@@ II.3 Probability distributions
In order to analyze the stochastic states encountered with positive $\Lambda$, we now consider the probability distribution $q^{(k)}(\varphi)$ arising from an initial $q^{(0)}(\varphi)$. The evolution of this $q(\varphi)$ can be directly inferred from that of individual phases $\varphi$ in Eq. (1).
Prior to the perturbation, the phase distribution remains unchanged up to rotation: $q^{(0)}(\varphi,t)=q^{(0)}(\varphi-t/T)$ (see illustration in SI Appendix, Fig. S1). We apply the same forcing (phase map $\psi$) after waiting times $\{\beta^{(k)}\}$ common to all oscillators. After the $k$’th kick, the ensemble will adopt the distribution $q^{(k)}(\varphi)$ according to
@@
For $\Lambda<0$, synchronization manifests in a very narrow, unimodal (“single-peaked”) $q^{(k)}(\varphi)$’s which continually become narrower on average. In contrast, for $\Lambda>0$ one finds that $q^{(k)}(\varphi)$ continues to change erratically with each kick, with no ultimate convergence to narrow distributions. In Fig. 2, we show the typical evolution of two distributions under $\Lambda<0$ and $\Lambda>0$.
@@
Figure 2: A sequence of two, initially-different phase distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ (red) and $q^{(k)}_{\mathrm{B}}(\varphi)$ (blue) subjected to common noise in a typical realization of Eq. (1) and Eq. (2) with either $\Lambda=-0.141$ ($A=7.32$) or $\Lambda=0.141$ ($A=9.32$), as indicated. The initial distributions are uniform, $\varphi\in[0.0,0.05)$ and $\varphi\in[0.5,0.55)$, respectively. The distributions are shown on a circle so the periodicity $\mathrm{mod}\,1$ of the phase circle is apparent. The radial axis shows the distributions’ values $q(\varphi)$ on a log scale, where $q(\varphi)=1$ for the inner full circle and as indicated for the outer dotted circles. The kick numbers $k$ are shown inside the inner circle. The $\Lambda<0$ dynamics are synchronizing, so their distributions become centered around the same phase value and their width decreases over time on average. For $\Lambda>0$, the distributions evolve erratically. However, remarkably, the red and blue points co-locate; they converge unto the same set of sharply-multimodal distributions under common noise — this property enables the effective synchronization demonstrated below. The distributions are sampled by tracking $N=500$ initial phases via the nearest-neighbor distances; see Eq. (8). The entropies of the two distributions and the KLDs among them are shown in Fig. 3 for $\Lambda=0.141$ and SI Appendix, Fig. S6 for $\Lambda=-0.141$. A detailed view of the blue distribution at $k=100$ for $\Lambda=0.141$ is shown in SI Appendix, Fig. S7.
We use Eq. (4) to follow the evolution of $q^{(k)}_{\mathrm{A}}(\varphi)$ and $q^{(k)}_{\mathrm{B}}(\varphi)$ for agent $\mathrm{A}$} and $\mathrm{B}$}. They start from different initial phase distributions $q^{(0)}_{\mathrm{A}}(\varphi)$ and $q^{(0)}_{\mathrm{B}}(\varphi)$, but both are subjected to the same phase map and sequence of kick timings $\{\beta^{(k)}\}$.
In Sec. III, we show that after sufficiently many kicks, the two distributions become operationally equivalent. This means that the same information is shared among the two agents, as both oscillators eventually follow the same statistics. To characterize this equivalence, we use the Kullback-Leibler divergence (KLD) [33] — a natural means of quantifying the divergence of the distribution $q_{\mathrm{A}}(\varphi)$ from $q_{\mathrm{B}}(\varphi)$,
@@
The convolution of Eq. (4) poses numerous issues: First, it cannot be carried out analytically for an arbitrary $q^{(k-1)}(\varphi)$. Second, each extremum in $\psi(\varphi)$, denoted $\varphi=\varphi_{\mathrm{m}}$, imparts an integrable singularity on $q^{(k)}(\varphi)$ of the form $\sim|\varphi-\psi(\varphi_{\mathrm{m}})|^{-1/2}$; these keep accumulating with each kick. Further, determining $q(\varphi)$ in practice requires sampling it with a finite number $N$ of samples. Thus, we sample the chosen initial distribution $q^{(0)}(\varphi)$ by $N$ samples $\{\varphi^{(0)}_{1},\varphi^{(0)}_{2},...\varphi^{(0)}_{N}\}$ drawn from $q^{(0)}(\varphi)$. Each of these $N$ samples changes under a kick via the phase map, according to Eq. (1). By repeatedly applying Eq. (1), we obtain the $\{\varphi^{(k)}_{1},\varphi^{(k)}_{2},...\varphi^{(k)}_{N}\}$, corresponding to a sample of $q^{(k)}(\varphi)$. We may then use standard tools to estimate $S^{(k)}$ [34] and $D^{(k)}(A||B)$ [35] in terms of the $\{\varphi^{(k)}_{n}\}$, as described in Appendix A.1. (The numerical uncertainty owing to discrete sampling in our simulations was of order $0.2$ for $S$ and $0.05$ for $D$ with $N=500$.) As we show in this work, the distributions of Fig. 2, obtained by such a discrete sampling, carry sufficient information about a synchronized phase notwithstanding the above singularities.

### v3

@@ Abstract
Two remote agents with synchronized clocks may use them to act in concert and communicate. This necessitates some means of creating and maintaining synchrony. One method, not requiring any direct interaction between the agents, is to expose them to a common, environmental, stochastic forcing. This “noise-induced synchronization” only occurs under sufficiently mild forcing; stronger forcing disrupts synchronization. We investigate the regime of strong noise, where the clocks’ relative phases evolve chaotically. Using a simple realization of disruptive noise, we demonstrate effective synchronization. First, although the relative phases of the two clocks varied erratically, we confirm that they became statistically independent of initial conditions and hence equivalent after a well-defined timescale. Second, we show that an agent can estimate an effective phase that closely agrees with the other’s phase. Thus, synchronization is practically attainable beyond the regime of conventional noise-induced synchronization.
@@ I Introduction
One such means is noise-induced synchronization, occurring when identical noninteracting oscillators are subjected to a common, nonlinear, stochastic perturbation from their environment [12, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. It is distinct from synchronization due to interactions between the clocks and from phase locking unto an external clock. Instead, synchronization is enabled via the clocks’ shared ambient noise. As such, noise-induced synchronization is the only means of maintaining synchronization between noninteracting clocks in a generic environment. When the noise is sufficiently mild, it can bring the two oscillators into synchrony over time. This has been well characterized experimentally in a range of physical contexts, such as electric circuits [21], quantum optics signaling [25], superconducting qubit chains [26, 27], and gene regulatory networks [28]. Synchronization can enable time-based communication between a sender and a receiver. If the sender emits an impulse when its clock is at a particular phase, the receiver can infer this phase by noting its own clock’s phase at the moment of the impulse.
@@
There is no guarantee that two oscillators will synchronize under a given noise. In general, synchronization requires the strength of the noise to lie below a threshold [12]. Beyond this threshold, the relative phase of the two oscillators devolves into chaos [18, 24]. In the latter case, synchronization is lost and with it the means of time-based communication outlined above. Nevertheless, we show below that a form of synchronization persists in the statistical properties of these erratically evolving phases. It permits each agent to define an effective phase that agrees with that of other agents to a quantitatively-predictable extent, thus enabling them to behave as though synchronized. We will refer to this property as effective synchronization.
@@
We demonstrate this behavior using a simplified type of noise, consisting of randomly timed impulses. For example, this may represent stochastically timed electric shocks delivered to a collection of firing neurons or injections of a chemical affecting the cell cycle [3] into a cell culture. We study the effect of these perturbations on a distribution of phases. Successive impulses create a sequence of phase distributions that in general depend on the initial distribution and the timings of the impulses. In below-threshold noise-induced synchronization, these distributions become narrow and unimodal, concentrated at a single phase. For noises above the threshold, the distributions remain spread over the entire phase circle [24].
@@
The above-threshold effective synchronization we report here arises by virtue of two properties: First, though the distributions change markedly with each impulse, we find that they become independent of the initial state after a finite and predictable number of impulses. This means that two initially different oscillator ensembles become statistically equivalent when compared at a given instant [29, 30, 31, 32]. Second, the distributions can be tightly bunched into a few sharp peaks. Indeed, when the threshold of synchronization is approached from above, the distributions show a high degree of order, since their average entropy becomes arbitrarily negative [24].
@@ II.1 Impulsive perturbations as phase maps
The effect of a kick on the clock can be compactly expressed via a phase-reduction prescription [22]: The oscillator traces a periodic loop in its dynamical manifold with a period $T$. We may label the points on this loop by a “phase position” $\varphi$ ranging from 0 to $1$, with a designated point assigned as the phase origin $\varphi\equiv 0$. Following Refs. [18, 22], we label other points to have phase values proportional to the time $\Delta t$ required for the oscillator to reach them from the phase origin: $\varphi\equiv\Delta t/T$. [^1] Unlike Refs. [18, 22], we do not include degrading noise, different for each oscillator. This eliminates the need to treat random excursions from the limit cycle via additional “radial” coordinates. We discuss the effect of degrading noise in Sec. V. (This construction is equivalent to finding the action-angle variables in Hamiltonian mechanics [33].) Thus, the phase of an oscillator that started at $\varphi^{(0)}$ advances at a constant speed, $\varphi^{(0)}+t/T$, where $\mathrm{mod}\,1$ is implied in any algebraic operation over the phases.
@@
A unique phase map $\psi(\varphi)$ can be determined for a given pairing of a dynamical system and a specified kick, as illustrated in Appendix A. In biological contexts, $\psi(\varphi)-\varphi$ is often called the phase-response curve, quantifying the response of circadian clocks [2, 34] and neural networks [35] to environmental changes. In these cases, the clock may be a living organism or periodically firing neurons, and the perturbations can be sporadic light impulses or shot noise from sensory neurons.
We study the case where this deterministic perturbation $\psi(\varphi)$ is applied at random times. [^2] We assume that the waiting times in between kicks are sufficiently long so the oscillators have relaxed back to the limit cycle. Thus to arbitrary accuracy, the complete physical state of each oscillator may be specified by its phase along this cycle. For the case of possibly insufficiently long times, see Ref. [36]. For simplicity, we may capture the effect of a $k$’th kick by observing the oscillators at an integer number of cycles after the latest kick. The unkicked position of the phase origin has continued its periodic motion and thus returns to the phase origin at such a moment. Likewise, the same phase is observed for the kicked oscillator after each completed cycle; we denote this phase as $\varphi^{(k)}$. Thus, the phase-map shift $\psi(\varphi^{\prime})$ from the unkicked phase origin is simply $\varphi^{(k)}$. To complete the iteration, we must express the phase position at the moment of the $k$th kick $\varphi^{\prime}$ in terms of the previous kick’s phase $\varphi^{(k-1)}$. During the random waiting time between the k1k-1’st and the $k$’th kick, the oscillator has gained a phase that we denote by $\beta^{(k)}$. Thus, $\varphi^{\prime}=\varphi^{(k-1)}+\beta^{(k)}$. Combining, we infer the effect of a single iteration,
@@
Immediately after a kick, this formula predicts the net phase shift that will be incurred after the oscillator has relaxed to its limit cycle. Due to the $\mathrm{mod}\,1$ constraint, only the fractional part of the waiting time in between kicks matters. For simplicity, we assume that $\{\beta^{(1)},\beta^{(2)},\ldots\}$ are identical independent uniformly-distributed random phases, $\beta^{(k)}\in[0,1)$. Equation (1), therefore, amounts to a Markov dynamics in the discrete “time” $k$, controlled by the phase map $\psi(\varphi)$ and the drawn $\beta^{(k)}$’s. Below, we compare the fates of two ensembles of oscillators exposed to the same sequence of $\beta^{(k)}$’s. [^3] We consider synchronization of distinct initial phase distributions under the influence of a shared forcing, as opposed to identical initial phase distribution evolving under different forcings. We average over $\{\beta^{(k)}\}$’s only after computing differences between the two agents obtained with fixed $\{\beta^{(k)}\}$’s.
@@
where the “gain parameter” $A$ determines the strength of the perturbations; when $A=0$, $\psi(\varphi)=\varphi$, so the kick has no effect. For completeness, we have repeated the paper’s quantitative analyses also for the quintic phase map, $\psi(\varphi)=\varphi+A^{\prime}\varphi(1-\varphi)(\Psi_{1}-\varphi)(\Psi_{2}-\varphi)(C-\varphi)$, where $C=(1-\Psi_{1})(1-\Psi_{2})/[\Psi_{1}\Psi_{2}+(1-\Psi_{1})(1-\Psi_{2})]$ is such that $(d\psi/d\varphi)|_{\varphi=1}=(d\psi/d\varphi)|_{\varphi=0}$. We will show the results for the cubic phase maps; we observed that the same effective synchronization occurs for the quintic ones.
@@ II.2 Lyapunov exponent and synchronization
With maps of this form, strong perturbations will typically be characterized by an increased gain parameter $A$. Once perturbed, the quantity $|d\psi/d\varphi|$ measures the degree of spreading of a local interval around $\varphi$. This “spreading factor” therefore is closely related to the (in)ability of the phase map to synchronize adjacent phases. Hence, a measure of effective strength of the kicks can be formulated by averaging this spreading factor. Accordingly, we define $\Lambda$ — the average of the log-spreading factor, [^4] The integration measure is uniform since the waiting times are distributed uniformly $\beta\in[0,1)$, so every oscillator is likely to experience a kick with equal probability everywhere along the phase circle.
@@ II.3 Probability distributions
In order to analyze the stochastic states encountered with positive $\Lambda$, we consider the probability distribution $q^{(k)}(\varphi)$ arising from an initial $q^{(0)}(\varphi)$. The evolution of this $q(\varphi)$ can be directly inferred from that of individual phases $\varphi$ in Eq. (1).
Prior to the perturbation, the phase distribution remains unchanged up to rotation: $q^{(0)}(\varphi,t)=q^{(0)}(\varphi-t/T)$ (see illustration in Appendix F). We apply the same forcing (phase map $\psi$) after waiting times $\{\beta^{(k)}\}$ common to all oscillators. After the $k$’th kick, the ensemble will adopt the distribution $q^{(k)}(\varphi)$ according to
@@
For $\Lambda<0$, synchronization manifests in a very narrow, unimodal (“single-peaked”) $q^{(k)}(\varphi)$’s which continually become narrower on average. In contrast, for $\Lambda>0$ one finds that $q^{(k)}(\varphi)$ continues to change erratically with each kick, with no ultimate convergence to narrow distributions.
@@
Figure 2: A sequence of two, initially-different phase distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ (red) and $q^{(k)}_{\mathrm{B}}(\varphi)$ (blue) subjected to common noise in a typical realization of Eq. (1) and Eq. (2) with either $\Lambda=-0.141$ ($A=7.32$) or $\Lambda=0.141$ ($A=9.32$), as indicated. The initial distributions are uniform, $\varphi\in[0.0,0.05)$ and $\varphi\in[0.5,0.55)$, respectively. The distributions are shown on a circle so the periodicity $\mathrm{mod}\,1$ of the phase circle is apparent. The radial axis shows the distributions’ values $q(\varphi)$ on a log scale, where $q(\varphi)=1$ for the inner full circle and as indicated for the outer dotted circles. The kick numbers $k$ are shown inside the inner circle. The $\Lambda<0$ dynamics are synchronizing, so their distributions become centered around the same phase value and their width decreases over time on average. For $\Lambda>0$, the distributions evolve erratically. However, remarkably, the red and blue points co-locate; they converge unto the same set of sharply-multimodal distributions under common noise — this property enables the effective synchronization demonstrated below. The distributions are sampled by tracking $N=500$ initial phases via the nearest-neighbor distances; see Eq. (9b). The entropies of the two distributions and the KLDs among them are shown in Fig. 3 for $\Lambda=0.141$ and Appendix E for $\Lambda=-0.141$. A detailed view of the blue distribution at $k=100$ for $\Lambda=0.141$ is shown in Appendix J.
We now compare the experience of two independent agents, $\mathrm{A}$} and $\mathrm{B}$}. Each has a copy of the clock and a knowledge of its phase map $\psi(\varphi)$. The two clocks have arbitrary phases. As time goes by, the two agents’ clocks experience a set of standard kicks at identical but random times. Each agent notes its oscillator’s phase at the time of kick $k$ (this phase is $\varphi^{\prime}$ in Fig. 1) from which $\beta^{(k)}$ can be calculated as we outlined in Sec. II.1. Knowing $\beta^{(k)}$, the agent can infer the phase shift incurred by any other clock $i$ whose phase was $\varphi_{i}^{(k-1)}$ before kick $k$, using Eq. (1). Indeed, it can infer the phases of $N$ clocks whose initial phases $\varphi_{i}^{(0)}$ were chosen randomly. It suffices to update the $i$’th phase at each successive kick. If agent $\mathrm{A}$} contains $N$ numerical registers that can each store a number from 0 to $1$, it can update each register with every kick and thus have $N$ samples of the phase after $k$ kicks. As described below, the agent can then use these samples to obtain an estimate distribution $q_{\mathrm{A}}^{(k)}(\varphi)$ of Eq. (4). Proceeding similarly, agent $\mathrm{B}$} can obtain an independent estimate $q_{\mathrm{B}}^{(k)}$. In Fig. 2, we show the typical evolution of two distributions under common noise with $\Lambda<0$ and $\Lambda>0$.
In Sec. III, we find that after a finite number of kicks, the two distributions become operationally equivalent. This means that the same information is shared among the two agents, as both oscillators eventually follow the same statistics. To characterize this equivalence, we use the Kullback-Leibler divergence (KLD) [37] — a natural means of quantifying the divergence of the distribution $q_{\mathrm{A}}(\varphi)$ from $q_{\mathrm{B}}(\varphi)$,
@@
The convolution of Eq. (4) poses numerous issues: First, it cannot be carried out analytically for an arbitrary $q^{(k-1)}(\varphi)$. Second, each extremum in $\psi(\varphi)$, denoted $\varphi=\varphi_{\mathrm{m}}$, imparts an integrable singularity on $q^{(k)}(\varphi)$ of the form $\sim|\varphi-\psi(\varphi_{\mathrm{m}})|^{-1/2}$; these keep accumulating with each kick. Further, determining $q(\varphi)$ in practice requires sampling it with a finite number $N$ of samples. Thus, we sample the chosen initial distribution $q^{(0)}(\varphi)$ by $N$ samples $\{\varphi^{(0)}_{1},\varphi^{(0)}_{2},...\varphi^{(0)}_{N}\}$ drawn from $q^{(0)}(\varphi)$. Each of these $N$ samples changes under a kick via the phase map, according to Eq. (1). By repeatedly applying Eq. (1), we obtain the $\{\varphi^{(k)}_{1},\varphi^{(k)}_{2},...\varphi^{(k)}_{N}\}$, corresponding to a sample of $q^{(k)}(\varphi)$. We may then use standard tools to estimate $S^{(k)}$ [38] and $D^{(k)}(A||B)$ [39] in terms of the $\{\varphi^{(k)}_{n}\}$, as described in Appendix B. (The numerical uncertainty owing to discrete sampling in our simulations was of order $0.2$ for $S$ and $0.05$ for $D$ with $N=500$.) As we show in this work, the distributions of Fig. 2, obtained by such a discrete sampling, carry sufficient information about a synchronized phase notwithstanding the above singularities.